# Perspective-Shifted Neuro-Symbolic World Models: A Framework for Socially-Aware Robot Navigation

### v1

@@ I INTRODUCTION & BACKGROUND
In particular, we seek to equip agents with the capability to reason, infer, and track others’ beliefs in social navigation tasks. Under a sequential decision-making framework, egocentric navigation can naturally be represented as an MDP. However, social navigation additionally requires reasoning about the hidden beliefs of others, inherently leading to a POMDP, where agents lack direct access to others’ mental states. Drawing inspiration from cognitive science studies on Theory of Mind and mental simulation [18], we instantiate these concepts within a neuro-symbolic MBRL architecture, where belief tracking informs decision-making and adaptation to social interactions. Theory of Mind grants individuals the ability to model their own and others’ mental states, fostering effective (and affective) collaboration without explicit communication or prior planning. Epistemic Planning provides a complementary logic-based framework enabling agents to explicitly reason about and manage the belief states of themselves and others during planning and execution [19]. We explore how Theory of Mind and Epistemic Planning can manifest in an intelligent agent architecture via mental simulation with learned representations.
@@ II PRELIMINARIES
Locality arises by focusing on a subset of factors $X=\{X_{1},\dots,X_{j}\}\subseteq S$ with $j\leq k$ that is relevant to a particular agent or sub-problem, with local state space $\mathcal{X}$ defined by $\mathcal{X}\;=\;\times_{i=1}^{j}\mathrm{dom}(X_{i})\;=\;\{\,\langle x_{1},\dots,x_{j}\rangle\mid x_{i}\in\mathrm{dom}(X_{i})\}$. Naively, one might attempt to track $x=\langle x_{1},\dots,x_{j}\rangle\in\mathcal{X}$ at each timestep by simply marginalizing out all other variables in $S$, but that is generally infeasible in large domains. Instead, IBA identifies a smaller set of non-local variables (influence sources) $u\in U\subseteq S\setminus x$ that directly affect $x$. This yields a local transition function $\dot{T}(x^{\prime}\mid x,u,a)$, a local observation function $\dot{O}(o\mid x,a)$, local reward function $\dot{R}(x,a,u)$ and an influence distribution $I(u\mid l)$, where $l$ denotes the action-local-state history needed to infer the current value of $u$. The action-local-state history up to time $t$ is $l_{t}=\langle x_{1},a_{1},\dots,a_{t-1},x_{t}\rangle$, where $x_{t}$ and $a_{t}$ are the local state and action at time $t$, respectively. Thus, $I(u_{t}\mid l_{t})$ captures the distribution over possible influences $u$ at time $t$, given the relevant local states and actions so far.
@@ III-B Perspective-Shift Operator
In social navigation, estimating the beliefs of other agents often requires reasoning from their perspective. Equipped with a world model, an agent can simulate hypothetical trajectories and their associated belief updates without real interactions with the environment. Assuming other agents share similar sensing and motion capabilities, we posit that estimating others’ beliefs in social navigation naturally entails perspective-taking[28, 29]. Formally, we introduce the perspective-shift operator $\Phi$:
$\tilde{b}_{\alpha}=\Phi(b,\hat{x}_{\alpha},M)$
where $b$ is the agent’s current (factored) belief, $\hat{x}_{\alpha}$ is the locally observed state configuration of another agent $\alpha$, and $M$ is the learned neuro-symbolic world model.
Intuitively, the operator $\Phi$ leverages the local transition and observation models from $M$ to simulate how the world would appear if the agent occupied the state configuration represented by $\hat{x}_{\alpha}$. Since we lack a global coordinate frame, the simulation treats the other agent’s configuration as a local target state, performing belief updates through imagined navigation steps towards that configuration. The resulting belief $\tilde{b}_{\alpha}$ encodes the observable environment from $\alpha$’s viewpoint, accounting for occlusions and partial observations that may differ significantly from the agent’s original perspective.
The perspective-shift operator thus serves as a computational instantiation of Theory of Mind, providing a mechanism to estimate the influence distribution $I(u\mid l)$ in the IALM formulation. Although focused specifically on social navigation, we envision future work extending $\Phi$ to generalize to other social reasoning tasks.
@@ III-C Influence-Augmented Policy Learning
Interestingly, by directly providing structured belief states as inputs, our approach effectively approximates a value function defined over a belief-MDP [31], aligning naturally with the POMDP formulation. This differs fundamentally from traditional use of the DQN algorithm in model-free settings, which typically approximates values directly from raw or unstructured observations. By combining structured symbolic belief-state representations (from the neuro-symbolic world model) with value-based policy learning (DQN), our approach bridges the gap between symbolic belief-MDP formulations and scalable deep RL methods, while simultaneously reducing the computational complexity associated with belief updates. Although we chose DQN primarily due to its simplicity and ease of implementation, future work could explore more specialized belief-space methods or policy-based approaches that may leverage the belief-MDP analogy more practically or efficiently. Furthermore, the explicit modeling of beliefs presented here opens promising avenues for integrating our neuro-symbolic world model with traditional belief-space planners and symbolic reasoning methods.
@@ IV EXPERIMENTS & RESULTS
We selected person-following as a representative social navigation task that naturally arises in real-world settings, such as service mobile robots operating in home environments, hospitals, or retail spaces. In such scenarios, robots are frequently expected to navigate socially; following and assisting humans to destinations or providing supportive behaviors. Person-following encompass key challenges in social navigation: reasoning about human intent, predicting future trajectories, handling partial observability, and maintaining socially acceptable distances (proxemics) and behaviors.
@@ IV-A Environment & Task
In our task (see Fig. 1), the agent’s objective is to efficiently follow a person-of-interest (POI) to an unknown goal without a predefined map, while adhering to proxemics and avoiding collisions. At each timestep, the agent receives positive reward for perceiving and staying near the POI, and negative reward for collisions. An additional large positive reward is given when the POI reaches the goal, terminating the episode [^2] Episodes are truncated if the agent strays too far from POI or exceeds a maximum number of steps.. To better align with realistic social navigation constraints, future implementations could explicitly encourage proxemics aligned with human preferences or promote clear visibility to the human.
@@ IV-C Results
footnotetext: Code available at: github.com/alcedok/SocialNav_paper

### v2

@@ I INTRODUCTION & BACKGROUND
In particular, we seek to equip agents with the capability to reason, infer, and track others’ beliefs in social-navigation tasks. Under a sequential decision-making framework, egocentric navigation is naturally an MDP, but introducing hidden beliefs turns the problem into a POMDP where agents lack direct access to others’ mental states. Drawing inspiration from cognitive science studies on Theory of Mind (ToM) and mental simulation [18], we embed belief tracking in a neuro-symbolic MBRL architecture that adapts decisions to social context. ToM supplies the mental-state concept, and Epistemic Planning provides the logic for acting over beliefs [19]. Guided by computational ToM models that cast belief inference as hidden variables in POMDPs [20], we restrict this study to first-order beliefs; while we do not yet exploit the full DEL guarantees of epistemic planners, our approach trades that formal rigor for data-driven scalability, and the symbolic latent factors could later interface with such formalisms. We therefore illustrate one way in which ToM and Epistemic Planning can co-manifest in a neuro-symbolic agent via latent-space mental simulation.
@@ II PRELIMINARIES
Locality arises by focusing on a subset of factors $X=\{X_{1},\dots,X_{j}\}\subseteq S$ with $j\leq k$ that is relevant to a particular agent or sub-problem, with local state space $\mathcal{X}$ defined by $\mathcal{X}\;=\;\times_{i=1}^{j}\mathrm{dom}(X_{i})\;=\;\{\,\langle x_{1},\dots,x_{j}\rangle\mid x_{i}\in\mathrm{dom}(X_{i})\}$. Naively, one might attempt to track $x=\langle x_{1},\dots,x_{j}\rangle\in\mathcal{X}$ at each timestep by simply marginalizing out all other variables in $S$, but that is generally infeasible in large domains. Instead, IBA identifies a smaller set of non-local variables (influence sources) $u\in U\subseteq S\setminus X$. This yields a local transition function $\dot{T}(x^{\prime}\mid x,u,a)$, a local observation function $\dot{O}(o\mid x,a)$, a local reward function $\dot{R}(x,a,u)$ and an influence distribution $I(u\mid l)$, where $l$ denotes the action-local-state history needed to infer the current value of $u$. The action-local-state history up to time $t$ is $l_{t}=\langle x_{1},a_{1},\dots,a_{t-1},x_{t}\rangle$, where $x_{t}$ and $a_{t}$ are the local state and action at time $t$, respectively. Thus, $I(u_{t}\mid l_{t})$ captures the distribution over possible influences $u$ at time $t$, given the relevant local states and actions so far.
@@ III-B Perspective-Shift Operator
In social navigation, estimating the beliefs of other agents often requires reasoning about their perspective. Equipped with a world model, an agent can simulate hypothetical trajectories and their associated belief updates without real interactions with the environment. Assuming other agents share similar sensing and motion capabilities, we posit that estimating others’ beliefs in social navigation entails perspective-taking [29, 30]. Formally, we introduce the perspective-shift operator $\Phi$:
$\tilde{b}_{\alpha}=\Phi(b,\hat{x}_{\alpha},\mathcal{M})$
where $b$ is the agent’s current (factored) belief, $\hat{x}_{\alpha}$ is the locally observed state configuration of another agent $\alpha$, and $\mathcal{M}=(\dot{T},\dot{O},\dot{R})$ is the learned neuro-symbolic world model.
Intuitively, the operator $\Phi$ leverages the local transition and observation models from $\mathcal{M}$ to simulate how the world would appear if the agent occupied the state configuration represented by $\hat{x}_{\alpha}$. Since we lack a global coordinate frame, the simulation treats the other agent’s configuration $\hat{x}_{\alpha}$ as a local target state and updates beliefs through imagined navigation. The resulting belief $\tilde{b}_{\alpha}$ encodes the observable environment from $\alpha$’s viewpoint, accounting for occlusions and partial observations that may differ significantly from the agent’s original perspective. The perspective-shift operator thus serves as a computational instantiation of Theory of Mind, providing a mechanism to estimate the influence distribution $I(u\mid l)$ in the IALM formulation. Although this work targets social navigation, we envision applying $\Phi$ to broader social‐reasoning tasks, e.g., collaborative assembly, where a perspective-shift would mentally simulate others’ limb poses and the assembly step in progress.
@@ III-C Influence-Augmented Policy Learning
By directly providing structured belief states as inputs, our approach effectively approximates a value function defined over a belief-MDP [32], aligning with the POMDP formulation. This differs fundamentally from traditional uses of DQN in model-free settings, where values are learned from raw observations. Combining symbolic belief-state representations produced by the neuro-symbolic world model with value-based policy learning bridges classical belief-MDP ideas and scalable deep RL, while reducing the computational complexity associated with belief updates. We adopt DQN because its value-based nature matches our discrete action space and aligns with classical value-based POMDP solvers, enabling direct future comparisons. Finally, explicit belief modeling opens avenues for integrating our neuro-symbolic world model with traditional belief-space planners and symbolic reasoning methods.
@@ IV EXPERIMENTS & RESULTS
We selected person-following as a representative social navigation task that arises in real-world settings, such as service mobile robots operating in home environments, hospitals, or retail spaces. In such scenarios, robots are frequently expected to navigate socially; following and assisting humans to destinations or providing supportive behaviors. Person-following encompass key challenges in social navigation: reasoning about human intent, predicting future trajectories, handling partial observability, and maintaining socially acceptable distances (proxemics) and behaviors.
@@ IV-A Environment & Task
In our task (see Fig. 1), the agent’s objective is to efficiently follow a person-of-interest (POI) to an unknown goal without a predefined map, while adhering to proxemics and avoiding collisions. At each timestep, the agent receives positive reward for perceiving and staying near the POI, and negative reward for collisions. An additional large positive reward is given when the POI reaches the goal, terminating the episode [^2] Episodes are truncated if the agent strays too far from POI or exceeds a maximum number of steps. Full environment, task, and reward specifications are available at github.com/alcedok/SocialNav_ROMAN25.. To better align with realistic social navigation constraints, future implementations could explicitly encourage proxemics aligned with human preferences or promote clear visibility to the human.